You bet on alotteryga ne. Y ugetS2 w achance of OA and $10 ith a chance of О The expected val e r th game s$140.4.20. 10. svou are risk a e se and wiling the game. if you pay $12 as a riok-neutral investor, what probabilities you would be assigning to the outcome of $20 and $10, respectively (0.e, risk neutral probabilities? pay y O Probability of $20 outcome Probability of $10 outcome 40% 60% O Probablity of $20 outcome Probablity of $10 outcome 20 Probability of $20outcome probat dity of %10 outcome 80% 20%
Answer #1

A risk neutral player will accept risk and play games with expected net payoffs of zero.A risk neutral player will accept risk and play games with expected net payoffs of zero.

so if we pay $12 as a risk neutral investor, then the risk neutral probabilities will be

Probability of $20 outcome is 20% and probability of $10 outcome is 80%

so the expected value will equal to $12

Expected value = 20%(20) + 80%(10) = 0.2*20 + 0.8*10 = 12

So correct ans is b) Probability of $20 outcome is 20% and probability of $10 outcome is 80%